# SToRM: Supervised Token Reduction for Multi-modal LLMs toward efficient end-to-end autonomous driving

### v1

@@ Abstract
In autonomous driving, end-to-end (E2E) driving systems that predict control commands directly from sensor data achieved significant advancements. For safe autonomous driving in unexpected scenarios, one may additionally rely on human interventions such as natural language instructions. Using a multi-modal large language model (MLLM) in autonomous driving facilitates human–vehicle interactions, and may improve driving performances in unexpected scenarios. However, this approach requires substantial computational resources due to its reliance on an LLM and many visual tokens from sensor inputs, that are inherently limited in autonomous vehicles. Many MLLM studies have explored reducing the number of visual tokens, and many approaches tend to exhibit some end-task performance degradation compared to using all tokens. For efficient E2E driving while maintaining driving performance comparable to using all tokens, this paper proposes the first Supervised Token Reduction framework for Multi-modal LLMs (SToRM). The proposed SToRM framework consists of three key elements. First, we propose a lightweight importance predictor with short-term sliding windows that predicts the importance scores of visual tokens. Second, we propose a supervised learning approach for the importance predictor, that uses an auxiliary path to obtain pseudo-supervision signals from an all-token pass through the LLM. Third, guided by predicted importance scores, we propose an anchor–context merging module that partitions tokens into “anchors” and “context” tokens, then merges the latter into their most relevant anchors to reduce redundancy while minimizing information loss. Experiments with the LangAuto benchmark dataset show that the proposed SToRM outperforms state-of-the-art E2E driving MLLM under an equal reduced-token budget and maintains all-token performance while substantially reducing computational cost, by up to $30\times$.
@@ II Related works
To handle challenging scenarios in E2E autonomous driving, one may use an MLLM with both visual and language information, promoting interactions between human decisions and information from multiple sensors [15, 5, 17]. In general, such MLLMs suffer from a large number of visual tokens generated from raw sensor inputs that leads to quadratic growth in computation cost due to the numerous attention layers in an LLM backbone. To accelerate the inference speed of E2E driving MLLM, it is crucial to reduce the number of visual tokens.
@@ III-A Modality encoding
At each time step $t\in\mathbb{N}_{\geq T}$, we generate a set of visual tokens $\{{\mathbf{Z}}_{t^{\prime}}\in\mathbb{R}^{N\times D}:t^{\prime}=t-T+1,\ldots,t\}$ from multi-view RGB images and a BEV map over the most recent $T$ frames, using the vision encoder pretrained by [15], where $N$ and $D$ denote the number of visual tokens for each frame and the embedding dimension of each visual token, respectively. The vision encoder backbone [15] consists of two convolutional neural network (CNN) encoders and a fusion transformer. A single CNN encoder is shared across all views to extract features from RGB images, while the other CNN encoder processes the BEV map. A fusion transformer then generates visual tokens by fusing the features extracted from the two encoders. We transform a language instruction into text tokens using the SentencePiece tokenizer [10].
@@ III-B1 Proposed importance predictor
for $\tau=1,\ldots,T$, where $\widetilde{{\mathbf{Z}}}_{\tau}$ is given in (1), $\sigma(\cdot)$ denotes the GeLU activation [9], ${\mathbf{W}}_{2}\in\mathbb{R}^{N\times 2|\mathcal{W}|\cdot N}$, ${\mathbf{W}}_{1}\in\mathbb{R}^{2|\mathcal{W}|\cdot N\times|\mathcal{W}|\cdot N}$, and $\widetilde{{\mathbf{Z}}}_{\mathcal{W}(\tau)}$ is given in (2). Note that $\{{\mathbf{W}}_{1},{\mathbf{W}}_{2}\}$ is shared for all $\tau$. See an illustration of the proposed processes (2)–(3) in Fig. 3(b).
@@ IV Experimental results and discussion
DS RC IS TFLOPs Memory
LLaVA (7B) All-token LMDrive 3,000 $44.0\pm 2.2$ $56.5\pm 3.8$ ${\mathbf{0.82}}\pm 0.02$ 30.80 16.20
LMDrive 120 $36.2\pm 2.3$ $46.5\pm 4.3$ $0.81\pm 0.03$ 1.04 14.10
SToRM (ours) 120 ${\mathbf{44.2}}\pm 2.5$ ${\mathbf{56.8}}\pm 3.3$ ${\mathbf{0.82}}\pm 0.02$ 1.02 14.10
TinyLLaVA (1.5B) All-token LMDrive 1,800 $39.6\pm 2.1$ $48.7\pm 4.0$ $0.82\pm 0.03$ 2.65 3.70
LMDrive 120 $30.9\pm 2.6$ $39.5\pm 4.1$ $0.82\pm 0.01$ 0.18 3.30
SToRM (ours) 120 ${\mathbf{40.8}}\pm 1.9$ ${\mathbf{49.3}}\pm 3.5$ ${\mathbf{0.84}}\pm 0.01$ 0.16 3.30
@@ IV-A Experimental setups
2) Evaluation metrics: We evaluated driving performance using route completion (RC), infraction score (IS), and driving score (DS). RC is the fraction of the route completed; $\text{IS}\!\in\![0,\!1]$ quantifies agent-triggered infractions; and DS captures both progress and safety as $\text{RC}\!\times\!\text{IS}$, reported as a percentage. We report inference efficiency in terms of floating-point operations (FLOPs; $1\,\text{GFLOP}=10^{9}\,\text{FLOPs}$; $1\,\text{TFLOP}={10}^{12}\,\text{FLOPs}$) or peak memory usage in gigabytes (GB). We report averages over three runs for all metrics.
@@ IV-B Comparisons between SToRM and SOTA E2E driving MLLM with different LLM scales
Across LLM backbones of different scales, Table II shows that the proposed SToRM achieves very comparable driving performances with SOTA E2E driving MLLM [15] using all tokens, i.e., all-token LMDrive, with far lower computational cost. In particular, compared with all-token LMDrive, SToRM reduces FLOPs by approximately $30\times$ with the large LLM backbone and $16.6\times$ with the tiny backbone. Tables II-III show that under the same visual-token budget, SToRM outperforms SOTA model [15], indicating efficient token use and robust generalization across LLM sizes. Finally, Table II shows that compared with the SOTA model, LMDrive (using LLaVA + Q-Former token reduction) [15], SToRM with the tiny LLM backbone, TinyLLaVA, attains significantly better driving performance while substantially improving inference efficiency, specifically, $6.5\times$ reduction in FLOPs and $4.27\times$ less memory.
@@ IV-E Comparisons between different visual token reduction approaches in ACM module
This section studies different token reduction schemes in the proposed ACM module in §III-B2: i) only use top-$K$ anchor tokens from our importance-based token categorization without merging, ii) a “soft” alternative that merges context tokens into all anchors with weighted contributions, and iii) the proposed “hard” scheme in (9), which merges each context token into its single most relavant anchor. Table VI shows that the proposed hard-assignment merging scheme outperforms the top-$K$-only selection scheme and the soft, naïve attention-based merging alternative. This suggests that hard merging preserves informative anchor representations, whereas soft merging introduces over-smoothing by merging contributions from all context tokens.
@@ V Conclusion
E2E autonomous driving models process sensor inputs (e.g., RGB images from camera(s) and point clouds from LiDAR(s)) and can benefit from language guidance. MLLM-based designs enable this capability but are computationally heavy due to large LLM backbones and many multi-modal tokens. For resource-constrained vehicles, it is therefore critical to reduce compute while maintaining performance.
We proposed SToRM, the first supervised token reduction framework for MLLMs that can significantly reduce computational cost without degrading E2E driving performance. The proposed SToRM outperformed the SOTA E2E driving MLLM [15] in both driving performances and inference efficiencies on the LangAuto benchmark.
The key idea of SToRM is to leverage pseudo-supervision signals to guide importance-aware token reduction. The proposed supervised learning approach uses an auxiliary path that provides pseudo-supervision signals extracted from an LLM using all tokens to an importance score predictor. The proposed importance predictor is lightweight by capturing short-term spatio-temporal relations among visual tokens. The proposed ACM module reduces visual tokens by first categorizing them into “anchor” and “context” groups and then merging the context tokens into anchors. On the LangAuto benchmark, SToRM surpassed SOTA E2E driving MLLM under an equal reduced-token budget and matches the all-token setting while cutting computation by up to $30\times$.
As a step toward (near) real-time multimodal E2E driving, we will integrate SToRM with complementary model-compression techniques (e.g., knowledge distillation and quantization).

### v3

@@ Abstract
In autonomous driving, end-to-end (E2E) driving systems that predict control commands directly from sensor data achieved significant advancements. For safe autonomous driving in unexpected scenarios, one may additionally rely on human interventions such as natural language instructions. Using a multi-modal large language model (MLLM) in autonomous driving facilitates human–vehicle interactions, and may improve driving performances in unexpected scenarios. However, this approach requires substantial computational resources due to its reliance on an LLM and many visual tokens from sensor inputs, that are inherently limited in autonomous vehicles. Many MLLM studies have explored reducing the number of visual tokens, and many approaches tend to exhibit some end-task performance degradation compared to using all tokens. For efficient E2E driving while maintaining driving performance comparable to using all tokens, this paper proposes the first Supervised Token Reduction framework for Multi-modal LLMs (SToRM). The proposed SToRM framework consists of three key elements. First, we propose a lightweight importance predictor with short-term sliding windows that predicts the importance scores of visual tokens. Second, we propose a supervised learning approach for the importance predictor, that uses an auxiliary path to obtain pseudo-supervision signals from an all-token pass through the LLM. Third, guided by predicted importance scores, we propose an anchor–context merging module that partitions tokens into “anchors” and “context” tokens, then merges the latter into their most relevant anchors to reduce redundancy while minimizing information loss. Experiments with the LangAuto benchmark dataset show that the proposed SToRM outperforms state-of-the-art E2E driving MLLM under an equal reduced-token budget and maintains all-token performance while substantially reducing computational cost, by up to $30\times$, and enabling real-time E2E driving on a standard GPU.
@@ II Related works
To handle challenging scenarios in E2E autonomous driving, one may use an MLLM with both visual and language information, promoting interactions between human decisions and information from multiple sensors. In general, such MLLMs suffer from a large number of visual tokens generated from raw sensor inputs that leads to quadratic growth in computation cost due to the numerous attention layers in an LLM backbone. To accelerate the inference speed of E2E driving MLLM, it is crucial to reduce the number of visual tokens.
@@ III-A Modality encoding
At each time step $t\in\mathbb{N}_{\geq T}$, we generate a set of visual tokens $\{{\mathbf{Z}}_{t^{\prime}}\in\mathbb{R}^{N\times D}:t^{\prime}=t-T+1,\ldots,t\}$ from multi-view RGB images and a BEV map over the most recent $T$ frames, using the vision encoder pretrained by [12], where $N$ and $D$ denote the number of visual tokens for each frame and the embedding dimension of each visual token, respectively. The vision encoder backbone [12] consists of two convolutional neural network (CNN) encoders and a fusion transformer. A single CNN encoder is shared across all views to extract features from RGB images, while the other CNN encoder processes the BEV map. A fusion transformer then generates visual tokens by fusing the features extracted from the two encoders. We transform a language instruction into text tokens using the SentencePiece tokenizer.
@@ III-B1 Proposed importance predictor
for $\tau=1,\ldots,T$, where $\widetilde{{\mathbf{Z}}}_{\tau}$ is given in (1), $\sigma(\cdot)$ denotes the GeLU activation, ${\mathbf{W}}_{2}\in\mathbb{R}^{N\times 2|\mathcal{W}|\cdot N}$, ${\mathbf{W}}_{1}\in\mathbb{R}^{2|\mathcal{W}|\cdot N\times|\mathcal{W}|\cdot N}$, and $\widetilde{{\mathbf{Z}}}_{\mathcal{W}(\tau)}$ is given in (2). Note that $\{{\mathbf{W}}_{1},{\mathbf{W}}_{2}\}$ is shared for all $\tau$. See an illustration of the proposed processes (2)–(3) in Fig. 3(b).
@@ IV Experimental results and discussion
DS RC IS TFLOPs Memory FPS
LLaVA (7B) All-token LMDrive 3,000 $44.0\pm 2.2$ $56.5\pm 3.8$ ${\mathbf{0.82}}\pm 0.02$ 30.80 16.20 4
LMDrive 120 $36.2\pm 2.3$ $46.5\pm 4.3$ $0.81\pm 0.03$ 1.04 14.10 24
SToRM (ours) 120 ${\mathbf{44.2}}\pm 2.5$ ${\mathbf{56.8}}\pm 3.3$ ${\mathbf{0.82}}\pm 0.02$ 1.02 14.10 25
TinyLLaVA (1.5B) All-token LMDrive 1,800 $39.6\pm 2.1$ $48.7\pm 4.0$ $0.82\pm 0.03$ 2.65 3.70 6
LMDrive 120 $30.9\pm 2.6$ $39.5\pm 4.1$ $0.82\pm 0.01$ 0.18 3.30 34
SToRM (ours) 120 ${\mathbf{40.8}}\pm 1.9$ ${\mathbf{49.3}}\pm 3.5$ ${\mathbf{0.84}}\pm 0.01$ 0.16 3.30 36
@@ IV-A Experimental setups
2) Evaluation metrics: We evaluated driving performance using route completion (RC), infraction score (IS), and driving score (DS). RC is the fraction of the route completed; $\text{IS}\!\in\![0,\!1]$ quantifies agent-triggered infractions; and DS captures both progress and safety as $\text{RC}\!\times\!\text{IS}$, reported as a percentage. We report inference efficiency in terms of floating-point operations (FLOPs; $1\,\text{GFLOP}=10^{9}\,\text{FLOPs}$; $1\,\text{TFLOP}={10}^{12}\,\text{FLOPs}$), peak memory usage in gigabytes (GB) and frames per second (FPS), measured on a single NVIDIA RTX 4090 GPU. We report averages over three runs for all metrics.
@@ IV-B Comparisons between SToRM and SOTA E2E driving MLLM with different LLM scales
Across LLM backbones of different scales, Table II shows that the proposed SToRM achieves very comparable driving performances with SOTA E2E driving MLLM [12] using all tokens, i.e., all-token LMDrive, with far lower computational cost. In particular, compared with all-token LMDrive, SToRM reduces FLOPs by approximately $30\times$ with the large LLM backbone and $16.6\times$ with the tiny backbone. These reductions make real-time E2E driving inference feasible on a standard GPU, with over $25$ FPS. Tables II-III show that under the same visual-token budget, SToRM outperforms SOTA model [12], indicating efficient token use and robust generalization across LLM sizes. Finally, Table II shows that compared with the SOTA model, LMDrive (using LLaVA + Q-Former token reduction) [12], SToRM with the tiny LLM backbone, TinyLLaVA, attains significantly better driving performance while substantially improving inference efficiency, specifically, $6.5\times$ reduction in FLOPs, $4.27\times$ less memory, and $1.5\times$ higher FPS.
@@ IV-E Comparisons between different visual token reduction approaches in ACM module
This section studies different token reduction schemes in the proposed ACM module in §III-B2: i) only use top-$K$ anchor tokens from our importance-based token categorization without merging, ii) a “soft” alternative that merges context tokens into all anchors with weighted contributions, and iii) the proposed “hard” scheme in (9), which merges each context token into its single most relevant anchor. Table VI shows that the proposed hard-assignment merging scheme outperforms the top-$K$-only selection scheme and the soft, naïve attention-based merging alternative. This suggests that hard merging preserves informative anchor representations, whereas soft merging introduces over-smoothing by merging contributions from all context tokens.
@@ V Conclusion
E2E autonomous driving models process multi-modal sensor inputs and can benefit from language guidance. However, MLLM-based designs are computationally heavy due to large LLM backbones and many multi-modal tokens. It is critical to reduce compute for resource-constrained vehicles.
We proposed SToRM, the first supervised token reduction framework for MLLMs that can significantly reduce computational cost without degrading E2E driving performance. The proposed SToRM outperformed the SOTA E2E driving MLLM [12] in both driving performances and inference efficiencies on the LangAuto benchmark. The key idea of SToRM is to leverage pseudo-supervision signals to guide importance-aware token reduction. SToRM employs a lightweight importance predictor and an ACM module that merges less important tokens into anchor tokens for efficient processing.